OpenStudy (anonymous):

One number is 12 greater than another.The product of the numbers is 28.Find the numbers. Can somebody give me the mother fluttering answer to this god damn problem!!!!! There are two pairs of numbers. One pair both of which are positive and the other both numbers are negative

OpenStudy (lilg132):

im helping you learn how to solve this

OpenStudy (anonymous):

i dont fluttering get whatyour saying idk what to substitute

OpenStudy (saifoo.khan):

Cool language. that can ban u from this website.

OpenStudy (anonymous):

let a and b represent both numbers. The first statement translates to \[a - b = 12 \] And the second statement translates to \[a * b = 24\] You can apply substitution to solve this problem. I'd suggest solving the first equation for a and then substituting for a.

OpenStudy (lilg132):

a = 12 + b